the segments GA and BG are tangent to a circle at A and B, and AGB is a 48 degree angle. Given that GA = 12 cm, find the distance from G to the nearest point on the circle. : the segments GA and BG are tangent to a circle at A and B, and AGB is a 48 degree angle. Given that GA = 12 cm, find the distance from G to the nearest point on the circle.
Answer by Edwin McCravy(2087)

the segments GA and BG are tangent to a circle at A and B, and AGB is a 48 degree angle. Given that GA = 12 cm, find the distance from G to the nearest point on the circle.

drawing(300,300,-6,14,-10,10, circle(0,0,5.343), locate(13.3,.4,G),
 locate(6,4,12), locate(10,.6,'48°'),locate(2.2,6,A), locate(2.2,-5,B),
line(13.13563534,0,2.173,-4.881),
line(13.13563534,0,2.173,4.881) ) 

Label the center of the circle O. and 
draw radii to A and B. Let the radius be r

drawing(300,300,-6,14,-10,10, circle(0,0,5.343), locate(13.3,.4,G),
 locate(6,4,12), locate(10,.6,'48°'),locate(2.2,6,A), locate(2.2,-5,B),
line(13.13563534,0,2.173,-4.881), line(2.173,-4.881,0,0),line(2.173,4.881,0,0),
line(13.13563534,0,2.173,4.881), locate(-1,.6,O),locate(.5,3.1,r), locate(.5,-2,r)

 )

Next draw OG which bisects the 48° angle G into
two 24° angles. Let P be the point where OG intersects
the circle.  P is the nearest point on the circle to G, 
so GP is the distance we're looking for.

Plan: Calculate the radius OA and the hypotenuse OG using the 
upper right triangle using trig ratios. Then calculate OG. Then 
since OP is also a radius, we will subtract the radius OP from 
OG and get GP.
 
drawing(300,300,-6,14,-10,10, circle(0,0,5.343), locate(13.3,.4,G),
 locate(6,4,12), locate(2.2,6,A), locate(2.2,-5,B),
line(13.13563534,0,2.173,-4.881), line(2.173,-4.881,0,0),line(2.173,4.881,0,0),
line(13.13563534,0,2.173,4.881), locate(-1,.6,O), line(13.13563534,0,0,0),
locate(9,1.2,'24°'), locate(9,-.4,'24°'), locate(5.4,1.2,P),locate(.5,3.1,r), locate(.5,-2,r)

  )

In the right triangle AOG, radius AO is the side opposite
angle AGO which is 24°.  GA is the side adjacent to angle AGO.

So we use 

tangent=(opposite)/(adjacent)

tan(24)=r/(GA)

tan(24)=r/12

Put 1 under the tan(24)

tan(24)/1=r/12

Cross-multiply:

r=12tan(24)

Next we calculate OG:

---
OG is the hypotenuse, GA is the opposite side of 24°

So we use 

cosine=(adjacent)/(hypotenuse)

cos(24)=GA/OG)

cos(24)=12/OG

Put 1 under the cos(24)

cos(24)/1=12/OG

Cross-multiply:

OG=12cos(24)

So now we can find GP by

subtraction, since OP = r = 12tan(24)

GP = OG - OP = 12cos(24)-12tan(24)=5.61981268cm

Edwin